OpenStudy (anonymous):

a competitor is aiming to complete a 1500-m wheel chair race in less than 4 mins. after moving at a constant speed for exactly 3.5 mins, there were still 240 m to go. what must his acceleration be for the remaining distance if he were to finish the race on time?

OpenStudy (anonymous):

The competitor moved 1260m in 3.5min (210s), his speed before acceleration is: 1260m/210s = 6m/s To determine a, use the equation: x = v*t + 1/2*a*t² a = 2(x - v*t)/t² a = 2(240m - 6m/s*30s) / (30s)² = 0.133333333 m/s² Answer: Acceleration must be greater than 0.13 m/s²